OpenStudy (anonymous):

For each of the following questions, choose the correct answer. Anita and Joelle bowl together and their combined total score for one game was 425 points. Anita’s score was 40 less than twice Joelle’s. What were their scores? Write a system of equations to model the problem if x represents Joelle’s score and y represents Anita’s score. A. x + y = 425 y = 2x + 40 B. x + y = 425 y = 2x – 40 C. x + y = 40 y = 2x – 425 D. x – y = 425 y = 2x – 40 @Directrix

OpenStudy (anonymous):

brb @Directrix look at the question while i am away please

Directrix (directrix):

The first step is the assignment of variables.

Directrix (directrix):

x represents Joelle’s score y represents Anita’s score

Directrix (directrix):

x + y = ?

OpenStudy (anonymous):

x=y=425

OpenStudy (anonymous):

sorry if i don't reply back fast I am eating while working

Directrix (directrix):

Is that how this happened? --> x=y=425 x + y + ?

OpenStudy (anonymous):

sorry messed up x+y=425

Directrix (directrix):

To avoid confusion, first code the equations with the letters of the participants' names; Anita’s score was 40 less than twice Joelle’s.

Directrix (directrix):

Anita’s score was 40 less than twice Joelle’s. A = 2*J - 40

OpenStudy (anonymous):

B is my guess

Directrix (directrix):

We are not doing answers yet. We are doing equations.

OpenStudy (anonymous):

ok ok sorry

Directrix (directrix):

A = 2*J - 40 -------------- Is Anita x or y? ---?

OpenStudy (anonymous):

y

Directrix (directrix):

Yes, so the equation goes to y = 2x - 40 Great news: It is time to pick the correct option.

OpenStudy (anonymous):

B

Directrix (directrix):

Yes, that is what I think.

Directrix (directrix):

Close and post.

OpenStudy (anonymous):

k